40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 1 22:03:44 2014, Koji, Summary, ASC, New PRM ASC is running 
    Reply  Wed Apr 2 19:50:12 2014, Koji, Summary, ASC, New PRM ASC is running 
       Reply  Wed Apr 2 23:08:51 2014, Koji, Summary, ASC, New PRM ASC is running PRM_QPD.pdfQPD_spe.pdf
Message ID: 9779     Entry time: Wed Apr 2 23:08:51 2014     In reply to: 9777
Author: Koji 
Type: Summary 
Category: ASC 
Subject: New PRM ASC is running 

The new PRM ASC design


Pitch
FM1: zero at 0Hz, pole at 2000Hz, gain at 2000Hz = 2000

FM5: (boost)
zero: f: 0.5Hz q: 1  /  4Hz, q: 2 / f: 1Hz, q: 3
pole: f: 2Hz q: 3  / f: 2.7Hz, q: 2  / f: 1Hz, q: 15

FM9: (HF Roll-off)
pole: f: 40Hz q: 1/Sqrt(2) (2nd order butterworth)

Servo gain: -0.023

Yaw
FM1: zero at 0Hz, pole at 2000Hz, gain at 2000Hz = 2000

FM5: (boost)
zero: f: 0.5Hz q: 1  /  4Hz, q: 2 / f: 1.5Hz, q: 10
pole: f: 1.02Hz q: 10  / f: 3Hz, q: 5  / f: 2Hz, q: 6

FM9: (HF Roll-off)
pole: f: 40Hz q: 1/sqrt(2)
 
Servo gain: -0.027


The loop gains were adjusted to have the UGFs of 10Hz. The measured openloop transfer functions were compared with the model.
The transfer functions for yaw are well matched. However, the pitch ones don't. It seems that the pitch loop has extra low pass
which I can't locate. The possibility is the analog electronics of the pitch loop.


The effect of the control between 0.3Hz to 3Hz are well represented by the model. The attachment 2 shows the free running
angle fluctuation, the ones with the control engaged, and the estimated spectra. Indeed, the estimated spectra well represent
the measured angular spectra.

Attachment 1: PRM_QPD.pdf  83 kB  Uploaded Thu Apr 3 00:14:37 2014  | Hide | Hide all
PRM_QPD.pdf
Attachment 2: QPD_spe.pdf  133 kB  Uploaded Thu Apr 3 00:14:49 2014  | Hide | Hide all
QPD_spe.pdf
ELOG V3.1.3-